OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle whose legs are 5 and √2. Please someone help asap! best answer/medal

OpenStudy (anonymous):

\[a ^{2}+b ^{2}=c ^{2}\]a=5\[b=\sqrt{2}\] \[5^{2}=25\]\[\sqrt{2}^{2}=2\] \[25+2=27\] \[\sqrt{27}=5.19\]